Que-1 a) According to the Langmuir–Hinshelwood mechanism of surface- catalyzed reactions, the rate of reaction between A and B depends on the rate at which the adsorbed species meet. Write the rate law for the reaction according to this mechanism. (b) Find the limiting form of this rate law for the case where the partial

pressures of the reactants are low. What is the overall order in this case? (c) Could

t


1
Expert's answer
2020-06-15T14:00:54-0400

(a)

Reaction is given by -


"aA+bB\to product"

Rate law -

"Rate =k[A]^m[B]^n"

(b)


"Rate = k[\frac{P_A}{RT}]^m[\frac{P_B}{RT}]^n\\=k[RT]^{-(m+n)}= constant \ at \ temp.\ T"

(Since partial pressure is low)